# Phase diagram of compressible and paired states in the quarter-filled Landau level

### v1

@@ IV.1 Laughlin state
We compute the energy expectation value of the $\nu=\frac{1}{3}$ Laughlin state for Coulomb interactions in the 0-7LL with exact diagonalization and using the Monte Carlo approach discussed above. For the former, we use the DiagHam librariesN. Regnault and et al. to generate Laughlin states with up to $N_{e}=14$ particles based on Ref. Bernevig and Regnault, 2009. Using Monte Carlo, we compute $G_{k}$ of the Laughlin state for various system sizes topping at $N_{e}=60$ electrons with $1.5\times 10^{9}$ uncorrelated updates. Using Eq. (16) we compute $V_{k}^{\text{$n$LL}}$ and refine the harmonics with procedure Eq. (18) to obtain $\tilde{V}^{\text{$n$LL}}_{k}$. Finally, we compute the energy Eq. (7) with Monte Carlo from the density-density correlation function $G_{k}$ and plot it as a function of $N_{e}^{-1}$ in Fig. 4; see Appendix B for our convention of normalizing energies. At system sizes where Monte Carlo and exact diagonalization are both applicable, the energies obtained with either approach match up to statistical errors. (For these system sizes, the error is too small to be visible in the scale of Fig. 4.)
@@ IV.2 Composite Fermi liquid
Composite Fermi liquids are metallic states that compete with paired states at half and quarter filling. Trial wavefunctions for such states can be obtained in a planar geometry by attaching the appropriate number $2p$ of flux quanta to a Slater determinant of plane-wave states,Rezayi and Read [1994] i.e.,
@@
where $\Phi_{n}$ denotes the Slater determinant of $n$ filled Landau levels. The number of filled Landau levels depends on the particle number via $N_{e}=n_{F}(n_{F}+2Q)$. As usual in Monte Carlo studies of quantum Hall wavefunctions, we employ a modified form of projection into the lowest Landau level to enable efficient simulations (cf. Appendix C).
@@ V Paired states
Within composite fermion theory, the even-denominator states at $\nu=\frac{1}{2p}$ can be viewed as a superfluid of composite fermions with $2p$ flux quanta. Read and Green [2000] This description can be used to deduce many properties of the states, such as their topological order, edge structure, and physical observables. In particular, different pairing channels give rise to distinct topological orders. Consequently, determining the favored pairing channel is an important question to study both experimentally and numerically.
To obtain a trial wavefunction for paired states, one replaces the Slater determinant in Eq. (20) with a Pfaffian factor. For an arbitrary odd pairing channel $\ell=2r+1$, we take Yutushui and Mross [2020]
@@
For $\ell=1$, this wavefunction reduces to the celebrated Moore-Read Pfaffian Moore and Read [1991], which does not require any explicit Landau-level projection. The anti-Pfaffian and f-wave topological orders correspond to $\ell=-3$ and $\ell=3$, respectively. For these two phases, different trial states based on the ‘parton’ framework are also available. Wen [1991] Their wavefunctions can be generated from $SU(2)_{\pm 2}$ chiral algebra Henderson et al. [2024] and take the form
@@
where $\Phi_{-2}\equiv\Phi^{*}_{2}$en [1991] The wavefunctions $\Psi_{SU(2)_{\pm 2}}$ differ microscopically from $\Psi_{\ell=\pm 3}$ but describe the same topological orders. The energies of parton and composite-fermion-based wavefunctions can differ, even after thermodynamic extrapolation, and we thus include both in our study. At half filling, the state $\Psi_{SU(2)_{2}}$ cannot be efficiently projected to the lowest Landau level (cf. Appendix C), and we use this state only at quarter filling.
@@ V.1 Half-filled states
The half-filled Landau level is the most extensively investigated even denominator; see Ref. Ma et al., 2022 for a recent review. Extensive numerical studies and experiments in many platforms, e.g., GaAs or graphene, agree that 0LLs with Coulomb favor a compressible CFL at this filling, while 1LL favors incompressible states. The nature of the latter remains the subject of an ongoing debate. Its resolution requires confronting a special feature of the half-filled Landau level compared to other even-denominators: Particle-hole symmetry ${\cal C}:\nu\rightarrow 1-\nu$ relates different competing phases. They are consequently degenerate for any two-body interactions in a single Landau level and considerable effort has been devoted to distinguish them in the presence of Landau-level mixing.
The filling factor $\nu=\frac{1}{2}$ is well suited for numerical works with particle numbers $N_{e}=2,4,\ldots,20$ readily accessible. Numerical studies based on exact diagonalization have found that the CFL in the 0LL quickly gives way to Moore-Read or anti-Pfaffian state upon tuning $H_{1}$ezayi and Haldane [2000], Möller and Simon [2008], Wójs et al. [2010] More recent work suggests that modifying the Coulomb interactions of the third Landau level in monolayer graphene by $H_{1}$ or $H_{3}$ can induce a state with f-wave pairing. Kim et al. [2019], Sharma et al. [2023]
@@
In our numerical studies at half filling, we included trial states for the three main pairing channels in addition to the CFL. Thermal transport measurements in GaAs indicate a fourth topological order known as the particle-hole symmetric PfaffianSon [2015]. However, no trial state for this phase in the lowest Landau level is known.Mishmash et al. [2018], Yutushui and Mross [2020], Rezayi et al. [2021]
Coulomb energies.—Firstly, we compute the energies of the trial states for Coulomb interactions in 0LL and 1LL. Fig. 6 shows the energies of the various trial states for system sizes up to $N_{e}=56$ ($N_{e}=70$ for the Moore-Read state, where explicit projection is not required). The data for all but the smallest systems are well described by linear fits to $N_{e}^{-1}$ (cf. Appendix B for details). In the 0LL, CFL is favored over all pairing channels, in agreement with previous works. By contrast, in 1LL, the Moore-Read and anti-Pfaffian wavefunctions have lower energies than the CFL. The linear fits for the Moore-Read and the anti-Pfaffian trial states extrapolate to somewhat different thermodynamic values. This discrepancy does not pose any contradiction to particle-hole symmetry. The specific trial states we use for the Moore-Read or anti-Pfaffian phases are not exact hole-conjugate of each other. Instead, ${\cal C}\Psi^{\nu=1/2}_{\ell=-3(1)}$ should be viewed as an alternative trial state for the Moore-Read (anti-Pfaffian) phase, which is microscopically distinct from $\Psi^{\nu=1/2}_{\ell=1(-3)}$. As a nonuniversal quantity, the energy is determined by microscopic properties and not only by the topological phase. Henderson et al. [2023], Sharma et al. [2021]
Haldane pseudopotentials.—Each paired state directly competes with a CFL at the same shift for $N_{e}$ such that an integer number of composite-fermion levels are filled. To understand which interactions favor paired states, we compute the energy difference between the paired states and CFLs for the first three pseudopotentials, i.e., ${\cal H}_{L}\equiv H_{L}\hat{\cal P}_{L}$ with $L=1,3$ and $5$, which are shown in Fig. 7. We find that CFL has the lowest energy for ${\cal H}_{1}$, the Moore-Read or anti-Pfaffian for ${\cal H}_{3}$, and f-wave for ${\cal H}_{5}$. Knowing that CFL is favored in the lowest Landau level, these results imply that reducing $H_{1}$ or increasing $H_{3}$ will drive a transition into an incompressible state with the Moore-Read or anti-Pfaffian pairing. The relative increase of a third over the first pseudopotential is also the primary difference between 0LL and 1LL Coulomb interactions. The f-wave state can instead by targeted by the $L=5$ pseudopotential, which is naturally prominent in the third Landau level of monolayer graphene. Kim et al. [2019]
@@
This phase diagram confirms qualitative features of tuning $\delta H_{1}$ and $\delta H_{3}$ obtained from exact diagonalization,Rezayi and Haldane [2000], Wójs et al. [2010] i.e., positive $\delta H_{1}$ and negative $\delta H_{3}$ derive system into gapless CFL phase.
@@ V.2 Quarter-filled states
At quarter filling, the Hilbert space size at the same particle number is significantly larger than at half filling. Consequently, exact diagonalization studies are limited to relatively small system sizes with $N_{e}\leq 12$, and Monte Carlo simulations play an important role in their study. Motivated by the experimental observation of quarter-filled states in wide quantum wells, Luhman et al. [2008], Shabani et al. [2009] numerical studies focused on finite-width effects. Papić et al. [2009a, b] These works found that the simplest model of the finite-width potentials Zhang and Das Sarma [1986] can drive the transition from CFL to the incompressible Moore-Read or Halperin (553) state. Later, other paired states were studied at quarter filling, and the f-wave state was found to be energetically favored by a realistic charge distribution in a wide quantum well. Faugno et al. [2019], Sharma et al. [2024] In our work, we systematically study generic interactions via their trial-state energies. Our findings can be readily applied to other systems where paired states were observed, such as hole-doped GaAs Wang et al. [2022, 2023] and bilayer graphene. Kumar et al. [2024]
@@
as a function of $x$, see Fig. 2. For the anti-Pfaffian and f-wave pairings, where the CFL wavefunctions Eq. (20) are defined for $N_{e}=10$, we observe a transition at a positive value of $x$ for which the paired states exhibit higher overlap with the ground state. The largest overlaps with the exact-diagonalization ground states are significantly larger than the ones between paired states and CFLs, see Appendix D. Meanwhile, negative $x$ favors CFL, which is reflected in an increasing overlap. We note that the transitions for small systems occur closer to the Coulomb point than in the thermodynamic limit. This behavior is consistent with Fig. 10, which shows larger energy differences between trial states at smaller particle numbers.
@@ VI Discussion
In this work, we developed a method of computing energies of quantum Hall trial wavefunctions for large systems with generic interactions. We used it to extract the thermodynamic energies of the first three Haldane pseudopotentials for different pairing channels of even-denominator states and their direct competitors, the CFLs. This allows us to construct the phase diagrams, Figs. 1 and 8, based on the energies of available trial states. We corroborated the phase diagram at quarter filling with exact diagonalization results; see Fig. 2. The results for half filling are consistent with earlier studies, e.g., Ref. Rezayi and Haldane, 2000, Wójs et al., 2010. Still, we caution that our analysis does not allow for other competing phases, such as Wigner crystals, which frequently occur at low filling factors.
@@
A recent experiment Ref. Kumar et al., 2024 observed an incompressible state at $\nu=\frac{3}{4}$ in several $N=0$ orbitals of bilayer graphene. By contrast, compressible states were observed at $\nu=\frac{1}{4}$ in the same Landau levels, which indicates a significant breaking of particle-hole symmetry, i.e., Landau-level mixing. Additional interactions arising from Landau-level mixing can thus tip the balance oppositely at $\nu=\frac{1}{4}$ and $\nu=\frac{3}{4}$. According to the phase diagram in Fig. 1, such interactions are most likely to favor the anti-Pfaffian phase, which is the closest to the 0LL Coulomb point. This result is compatible with a satellite plateau that has been observed at the filling factor $\nu=\frac{19}{25}$, consistent with anti-Pfaffian topological order. Yutushui et al. [2024]
@@
Finally, we expect that one could generalize our approach to many-body interactions. Specifically, three-body pseudopotentials are commonly used in exact diagonalization studies of Landau level mixing effects. Computing their expectation values with Monte Carlo would enable studying the effects of Landau level mixing on larger systems.
@@ Appendix A Higher Landau levels
On a sphere with monopole strength $q^{\prime}=l-n$, the number of states in the $n$th Landau level is $12l+$ 1. Consequently, any two-body interactions in this Landau level are parametrized by $12l+$ 1 Haldane pseudopotentials $H^{\text{$n$LL}}_{L}$ (cf. the main text, where $12q+$ 1 pseudopotentials $H_{L}$ parameterize the lowest Landau level interactions). Given the harmonics $V_{k}$, these pseudopotentials can be computed Wooten [2013] as
$H^{\text{$n$LL}}_{L}=\sum_{k=0}^{2l}M_{L,k}(l,q^{\prime}=l-n)V_{k}`.$ (27)

### v3

@@ IV.1 Laughlin state
We compute the energy expectation value of the $\nu=\frac{1}{3}$ Laughlin state for Coulomb interactions in the 0-7LL with exact diagonalization and using the Monte Carlo approach discussed above. For the former, we use the DiagHam librariesN. Regnault and et al. to generate Laughlin states with up to $N_{e}=14$ particles based on Ref. Bernevig and Regnault, 2009. Using Monte Carlo, we compute $G_{k}$ of the Laughlin state for various system sizes topping at $N_{e}=60$ electrons with $1.5\times 10^{9}$ uncorrelated updates. Using Eq. (16) we compute $V_{k}^{\text{$n$LL}}$ and refine the harmonics with procedure Eq. (18) to obtain $\tilde{V}^{\text{$n$LL}}_{k}$. Finally, we compute the energy Eq. (7) with Monte Carlo from the density-density correlation function $G_{k}$ and plot it as a function of $N_{e}^{-1}$ in Fig. 4; see Appendix B for our convention of normalizing energies. At system sizes where Monte Carlo and exact diagonalization are both applicable, the energies obtained with either approach match up to statistical errors. (For these system sizes, the error is too small to be visible in the scale of Fig. 4.) The lowest Landau level energies and thermodynamic extrapolations reproduce values previously reported in the literature. Ciftja and Wexler (2003); Jain (2007)
@@ IV.2 Composite Fermi liquid
Composite Fermi liquids are metallic states that compete with paired states at half and quarter filling. Trial wavefunctions for such states can be obtained in a planar geometry by attaching the appropriate number $2p$ of flux quanta to a Slater determinant of plane-wave states,Rezayi and Read (2005) i.e.,
@@
where $\Phi_{n}$ denotes the Slater determinant of $n$ filled Landau levels. The number of filled Landau levels depends on the particle number via $N_{e}=n_{F}(n_{F}+2Q)$. As usual in Monte Carlo studies of quantum Hall wavefunctions, we employ a modified form of projection into the lowest Landau level to enable efficient simulations (cf. Appendix D).
@@ V Paired states
Within composite fermion theory, the even-denominator states at $\nu=\frac{1}{2p}$ can be viewed as a superfluid of composite fermions with $2p$ flux quanta. Read and Green (2000) This description can be used to deduce many properties of the states, such as their topological order, edge structure, and physical observables. In particular, different pairing channels give rise to distinct topological orders. Consequently, determining the favored pairing channel is an important question to study both experimentally and numerically.
To obtain a trial wavefunction for paired states, one replaces the Slater determinant in Eq. (20) with a Pfaffian factor. For an arbitrary odd pairing channel $\ell=2r+1$, we take Yutushui and Mross (2020)
@@
For $\ell=1$, this wavefunction reduces to the celebrated Moore-Read Pfaffian Moore and Read (1991), which does not require any explicit Landau-level projection. The anti-Pfaffian and f-wave topological orders correspond to $\ell=-3$ and $\ell=3$, respectively. For these two phases, different trial states based on the ‘parton’ framework are also available. Jain (1989b); Wen (1991) Their wavefunctions can be generated from $SU(2)_{\pm 2}$ chiral algebra Henderson et al. (2024) and take the form
@@
where $\Phi_{-2}\equiv\Phi^{*}_{2}$en (1991) The wavefunctions $\Psi_{SU(2)_{\pm 2}}$ differ microscopically from $\Psi_{\ell=\pm 3}$ but describe the same topological orders. Balram et al. (2018) The energies of parton and composite-fermion-based wavefunctions can differ, even after thermodynamic extrapolation, and we thus include both in our study. At half filling, the state $\Psi_{SU(2)_{2}}$ cannot be efficiently projected to the lowest Landau level (cf. Appendix D), and we use this state only at quarter filling.
@@ V.1 Half-filled states
The half-filled Landau level is the most extensively investigated even denominator; see Ref. Ma et al., 2024 for a recent review. Extensive numerical studies and experiments in many platforms, e.g., GaAs or graphene, agree that 0LLs with Coulomb favor a compressible CFL at this filling, while 1LL favors incompressible states. The nature of the latter remains the subject of an ongoing debate. Its resolution requires confronting a special feature of the half-filled Landau level compared to other even-denominators: Particle-hole symmetry ${\cal C}:\nu\rightarrow 1-\nu$ relates different competing phases. They are consequently degenerate for any two-body interactions in a single Landau level and considerable effort has been devoted to distinguish them in the presence of Landau-level mixing.
The filling factor $\nu=\frac{1}{2}$ is well suited for numerical works with particle numbers $N_{e}=2,4,\ldots,20$ readily accessible. Numerical studies based on exact diagonalization have found that the CFL in the 0LL quickly gives way to Moore-Read or anti-Pfaffian state upon tuning $H_{1}$ezayi and Haldane (2000); Möller and Simon (2008); Wójs et al. (2010) More recent work suggests that modifying the Coulomb interactions of the third Landau level in monolayer graphene by $H_{1}$ or $H_{3}$ can induce a state with f-wave pairing. Kim et al. (2019); Sharma et al. (2023)
@@
In our numerical studies at half filling, we included trial states for the three main pairing channels in addition to the CFL. Thermal transport measurements in GaAs indicate a fourth topological order known as the particle-hole symmetric PfaffianSon (2015). However, no trial state for this phase in the lowest Landau level is known.Mishmash et al. (2018); Yutushui and Mross (2020); Rezayi et al. (2021)
Coulomb energies.—Firstly, we compute the energies of the trial states for Coulomb interactions in 0LL and 1LL. Fig. 6 shows the energies of the various trial states for system sizes up to $N_{e}=56$ ($N_{e}=70$ for the Moore-Read state, where explicit projection is not required). The data for all but the smallest systems are well described by linear fits to $N_{e}^{-1}$ (cf. Appendix B for details). In the 0LL, CFL is favored over all pairing channels, in agreement with previous works. By contrast, in 1LL, the Moore-Read and anti-Pfaffian wavefunctions have lower energies than the CFL. The linear fits for the Moore-Read and the anti-Pfaffian trial states extrapolate to somewhat different thermodynamic values. This discrepancy does not pose any contradiction to particle-hole symmetry. The specific trial states we use for the Moore-Read or anti-Pfaffian phases are not exact hole-conjugate of each other. Instead, ${\cal C}\Psi^{\nu=1/2}_{\ell=-3(1)}$ should be viewed as an alternative trial state for the Moore-Read (anti-Pfaffian) phase, which is microscopically distinct from $\Psi^{\nu=1/2}_{\ell=1(-3)}$. As a nonuniversal quantity, the energy is determined by microscopic properties and not only by the topological phase. Henderson et al. (2023); Sharma et al. (2021)
Haldane pseudopotentials.—Each paired state directly competes with a CFL at the same shift for $N_{e}$ such that an integer number of composite-fermion levels are filled. To understand which interactions favor paired states, we compute the energy difference between the paired states and CFLs for the first three pseudopotentials, i.e., ${\cal H}_{L}\equiv H_{L}\hat{\cal P}_{L}$ with $L=1,3$ and $5$, which are shown in Fig. 7 (it becomes increasingly more complicated to reliably extract the energy for individual pseudopotentials with larger $L$; see Appendix C). We find that CFL has the lowest energy for ${\cal H}_{1}$, the Moore-Read or anti-Pfaffian for ${\cal H}_{3}$, and f-wave for ${\cal H}_{5}$. Knowing that CFL is favored in the lowest Landau level, these results imply that reducing $H_{1}$ or increasing $H_{3}$ will drive a transition into an incompressible state with the Moore-Read or anti-Pfaffian pairing. The relative increase of a third over the first pseudopotential is also the primary difference between 0LL and 1LL Coulomb interactions. The f-wave state can instead by targeted by the $L=5$ pseudopotential, which is naturally prominent in the third Landau level of monolayer graphene. Kim et al. (2019)
@@
This phase diagram confirms qualitative features of tuning $\delta H_{1}$ and $\delta H_{3}$ obtained from exact diagonalization,Rezayi and Haldane (2000); Wójs et al. (2010) i.e., positive $\delta H_{1}$ and negative $\delta H_{3}$ derive system into gapless CFL phase.
@@ V.2 Quarter-filled states
At quarter filling, the Hilbert space size at the same particle number is significantly larger than at half filling. Consequently, exact diagonalization studies are limited to relatively small system sizes with $N_{e}\leq 12$, and Monte Carlo simulations play an important role in their study. Motivated by the experimental observation of quarter-filled states in wide quantum wells, Luhman et al. (2008); Shabani et al. (2009) numerical studies focused on finite-width effects. Papić et al. (2009a, b) These works found that the simplest model of the finite-width potentials Zhang and Das Sarma (1986) can drive the transition from CFL to the incompressible Moore-Read or Halperin (553) state. Later, other paired states were studied at quarter filling, and the f-wave state was found to be energetically favored by a realistic charge distribution in a wide quantum well. Faugno et al. (2019); Sharma et al. (2024) In our work, we systematically study generic interactions via their trial-state energies. Our findings can be readily applied to other systems where paired states were observed, such as hole-doped GaAs Wang et al. (2022, 2023) and bilayer graphene. Kumar et al. (2024) In particular, the main results in Fig. 1 can be viewed as perturbations of the $N=0$ orbital of bilayer graphene since the interactions there are the same as in the 0LL of narrow GaAs quantum wells.
@@
as a function of $x$, see Fig. 2. For the anti-Pfaffian and f-wave pairings, where the CFL wavefunctions Eq. (20) are defined for $N_{e}=10$, we observe a transition at a positive value of $x$ for which the paired states exhibit higher overlap with the ground state. The largest overlaps with the exact-diagonalization ground states are significantly larger than the ones between paired states and CFLs, see Appendix E. Meanwhile, negative $x$ favors CFL, which is reflected in an increasing overlap. We note that the transitions for small systems occur closer to the Coulomb point than in the thermodynamic limit. This behavior is consistent with Fig. 10, which shows larger energy differences between trial states at smaller particle numbers.
@@ VI Discussion
In this work, we developed a method of computing energies of quantum Hall trial wavefunctions for large systems with generic interactions. We used it to extract the thermodynamic energies of the first three Haldane pseudopotentials for different pairing channels of even-denominator states and their direct competitors, the CFLs. This allows us to construct the phase diagrams, Figs. 1 and 8, based on the energies of available trial states. We corroborated the phase diagram at quarter filling with exact diagonalization results; see Fig. 2. The results for half filling are consistent with earlier studies, e.g., Ref. Rezayi and Haldane, 2000; Wójs et al., 2010. Still, we caution that our analysis does not allow for other competing phases, such as Wigner crystals, which frequently occur at low filling factors. Ultimately, the direct identification of pairing channel can be achieved through thermal transport experiments; Banerjee et al. (2018); Dutta et al. (2022a, b); Melcer et al. (2022); Paul et al. (2024); Melcer et al. (2024); Paul et al. (2024); Dutta et al. (2022c) see also other proposals. Manna et al. (2024); Yutushui et al. (2022); Park et al. (2024); Yutushui and Mross (2023); Yutushui et al. (2024b, 2024)
@@
A recent experiment Ref. Kumar et al., 2024 observed an incompressible state at $\nu=\frac{3}{4}$ in several $N=0$ orbitals of bilayer graphene. By contrast, compressible states were observed at $\nu=\frac{1}{4}$ in the same Landau levels, which indicates a significant breaking of particle-hole symmetry, i.e., Landau-level mixing. Additional interactions arising from Landau-level mixing can thus tip the balance oppositely at $\nu=\frac{1}{4}$ and $\nu=\frac{3}{4}$. According to the phase diagram in Fig. 1, such interactions are most likely to favor the anti-Pfaffian phase, which is the closest to the 0LL Coulomb point. This result is compatible with a satellite plateau that has been observed at the filling factor $\nu=\frac{19}{25}$, consistent with anti-Pfaffian topological order. Yutushui et al. (2024a)
@@
Finally, we expect that one could generalize our approach to many-body interactions. Specifically, three-body pseudopotentials are commonly used in exact diagonalization studies of Landau level mixing effects. Computing their expectation values with Monte Carlo would enable studying the effects of Landau level mixing on larger systems. We note, however, that Landau level mixing does not always break particle-hole symmetry, e.g., in the zeroth Landau level of graphene with electrons and holes excited symmetrically. In such cases, the three-body interactions vanish, and our method can be readily used to assess the effects of Landau level mixing on energies.
Note added—The results of Ref. Huang and Wu, 2024, which studied the $\nu=3/4$ filled states in bilayer graphene, are consistent with our finding. A recent study Ref. Nešković et al., 2024 relates pairing between composite fermions to interactions between their dipole moments.
@@ Acknowledgements.
DATA AVAILABILITY
Data used in the paper is publicly available in Zenodo at 10.5281/zenodo.14030395.
@@ Appendix A Higher Landau levels
On a sphere with monopole strength $q^{\prime}=l-n$, the number of states in the $n$th Landau level is $12l+$ 1. Consequently, any two-body interactions in this Landau level are parametrized by $12l+$ 1 Haldane pseudopotentials $H^{\text{$n$LL}}_{L}$ (cf. the main text, where $12q+$ 1 pseudopotentials $H_{L}$ parameterize the lowest Landau level interactions). Given the harmonics $V_{k}$, these pseudopotentials can be computed Wooten (2013) as
$H^{\text{$n$LL}}_{L}=\sum_{k=0}^{2l}M_{L,k}(l,q^{\prime}=l-n)V_{k}.$ (27)